# Quantum Entanglement Trees: Optimizing Quantized Matrix Quantization via Element Replacement and Residual Clustering

### v3

@@ 7 Conclusion
In conclusion, our research addresses critical challenges in the optimization of large language models (LLMs) through the development of a novel algorithm, HERA. The growing complexity and scale of LLMs necessitate innovative approaches to enhance storage and computational efficiency, particularly in vector database management, weight quantization, and key-value (k-v) cache optimization.
HERA’s hierarchical approach to segmenting, compressing, and reorganizing the matrix dataset has proven to be effective in significantly reducing quantization error. By considering the distribution and magnitude of parameters, our method achieves superior performance compared to traditional uniform quantization techniques. The experimental results from our prototype system, implemented in Python, demonstrate that HERA can reduce the quantization error to 12.3% of the original, maintaining the same compression ratio.
Future work will explore further refinements to HERA, including its application to other types of application and broader datasets. Additionally, investigating the integration of HERA with other optimization techniques may yield even greater enhancements in model performance and efficiency. The promising results of this study encourage continued research and development in the quest for more effective and scalable solutions for LLM compression and optimization.

### v4

@@ 6 Conclusion
Matrix quantization involves representing matrix elements in a more space-efficient form to reduce storage usage, with dequantization restoring the original matrix for practical applications. We frame the Quantization Error Minimization (QEM) problem as minimizing the distance between a matrix before and after quantization, constrained by the condition that the quantized matrix occupies the same memory space. Matrix quantization plays a crucial role in various domains, including Large Language Models (LLMs) weight quantization, vector databases, KV cache quantization, graph compression, and image compression. Recent advancements in LLMs, such as GPT-4 and BERT, have underscored the importance of matrix compression due to the large size of parameters and KV caches, which are stored as matrices.
To address the QEM problem, we propose Quantum Entanglement Trees (QET), which leverage the local orderliness of matrix elements through iterative element swapping to form a locally ordered matrix. This matrix is subsequently grouped and quantized by columns. To enhance the QET algorithm, we introduce two key optimizations: further quantizing residuals to reduce Mean Squared Error (MSE) and utilizing masking and batch processing techniques to accelerate the algorithm.
Our experimental results demonstrate that QET can effectively reduce MSE to 5.05%, 13.33%, and 11.89% of the best existing methods on the LLM dataset, K cache, and V cache, respectively. These findings highlight QET’s efficiency in reducing quantization error while maintaining compression effectiveness.
The contributions of this work include the formal abstraction of the QEM problem, the design of the QET algorithm, and the introduction of two optimizations that improve both accuracy and computational speed. Future work will focus on extending this algorithm to the quantization of large-scale model parameters and KV caches. This will allow us to further evaluate the algorithm’s effectiveness in addressing the challenges posed by large-scale models and real-time processing scenarios, thereby broadening its applicability and impact in practical machine learning tasks.